Circular No. 7626

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bp IN ANONYMOUS GALAXY M. Papenkova and W. D. Li, University of California at Berkeley

(UCB), report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 17.5) in an unfiltered image taken with the Katzman Automatic Imaging Telescope (KAIT) on May 15.4 UT. The new object was confirmed at about the same brightness on an unfiltered image taken with the Tenagra Observatory 0.5-m automatic telescope by M. Schwartz on May 16.2 UT. SN 2001bp is located at α = 16^{h}02^{m}09^{s}.25, δ = +36^{o}43′08′′.0 (equinox 2000.0), which is 3′′.5 east and 6′′.0 south of an anonymous galaxy. A KAIT image of the same field taken on May 8.5 UT showed nothing at this position (limiting magnitude ∼ 19.0). A finding chart can be found at

http://astron.berkeley.edu/~bait/2001/sn0515.gif.

A. V. Filippenko and R. Chornock, UCB, report that inspection of a CCD spectrum (range 320–1000 nm) obtained on May 16 UT with the Shane 3-m reflector at the Lick Observatory reveals that the object is a young type Ia supernova close to, or before, maximum brightness. Its redshift, measured from the supernova features and from narrow emission lines in the host galaxy, is ∼ 0.095.

SUPERNOVA 2001bg IN NGC 2608 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001bg, obtained by M. Calkins on May 15.23 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), confirms it to be a type Ia supernova near maximum light (cf. IAUC 7622). Adopting the NASA/IPAC Extragalactic Database recession velocity of 2135 km/s for the host galaxy, we find the supernova expansion velocity from Si II (rest 635.5 nm) to be ∼ 12 000 km/s, which differs from the velocity of 9700 km/s reported on IAUC 7622. There is a strong (EW ∼ 0.15 nm) Na I D absorption line at the velocity of the host galaxy and a weaker (EW ∼ 0.05 nm) Na I D absorption at zero velocity. The spectral feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 5 ± 2 days after maximum light.

SUPERNOVA 1995N IN MCG −02-38-017 B. E. Schaefer, University of Texas at Austin, reports on photometry

of a type IIn supernova obtained with the 2.1-m telescope at the McDonald Observatory: “SN 1995N has V = 21.1 ± 0.3 and R = 20.5 ± 0.2 in observations made on Apr. 19.3 UT, 2175 days after discovery (IAUC 6170). It has faded 1.3 mag in V over the last 770 days (IAUC 7141).”

2001 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden